OpenStudy (anonymous):

Find the x-coordinates of any relative extrema and inflection point(s) for the function f of x equals 9 times x raised to the one third power plus 9 halves times x raised to the four thirds power . You must justify your answer using an analysis of f ′(x) and f ′′(x).

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

\[f(x)=9x ^{\frac{ 1 }{ 3 }}+\frac{ 9 }{ 2 }x^\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

is that the given equation?

OpenStudy (anonymous):

@1019.jams do you know how to take the first and second derivatives of f(x)?

OpenStudy (anonymous):

yes :) & im not sure! no i dont think so can you please show me :)

OpenStudy (anonymous):

@1019.jams you have learned about taking derivatives using the "power rule"?

OpenStudy (anonymous):

if f(x) = x^n then f'(x) = nx^(n-1)